# Spectrum of random centrosymmetric matrices; CLT and Circular law

### v2

@@ Abstract.
Indrajit Jana’s research is partially supported by INSPIRE Fellowship DST/INSPIRE/04/2019/000015, Dept. of Science and Technology, Govt. of India.
@@ 2. Notations and layout
This article is organized as follows. We state the main theorem in Section 3. The proof of the CLT is shown in Section 6, and the variance calculation of the limiting normal distribution is delegated to Section 7. In addition, it was also argued that the limiting spectral distribution follows the circular law, which was demonstrated in Section 5.
Throughout this article, the identity matrix is denoted by $I$, and the counter-identity matrix is denoted by $J$, as defined in Definition 4.1. The dimensions of such matrices should be understood from the context. The eigenvalues of an $n\times n$ matrix $M$ are denoted by $\lambda_{1}(M),$ $\lambda_{2}(M),$ $\ldots$,… , $\lambda_{n}(M).$ The complex conjugate transpose of a matrix $Y$ is denoted by $Y{\operatorname{{}^{*}}}$, whereas $\bar{Y}$ stands for only the complex conjugate of $Y$ without taking transpose. The notation $\mathcal{R}_{z}(M)=(zI-M)^{-1}$ represents the resolvent of a square matrix $M$. The notations $\zeta_{n}\stackrel{{\scriptstyle d}}{{\to}}\zeta$, $\zeta_{n}\stackrel{{\scriptstyle p}}{{\to}}\zeta$, and $\zeta_{n}\stackrel{{\scriptstyle a.s.}}{{\to}}\zeta$ are used to indicate that the sequence of random variables $\{\zeta_{n}\}_{n}$ converges to $\zeta$ in distribution, in probability, and almost surely, respectively.
@@ 3. Main result
Figure 1. Symmetry pattern of a centrosymmetric 5$\times$5 matrix.
@@
Our main theorem is stated as follows.
@@
[…]
In particular, if $f$ is a polynomial given by $P_{d}(z)=\sum_{k=0}^{d}a_{k}z^{k}$, then $\sigma_{p}^{2}=\sum_{k=1}^{d}2k|a_{k}|^{2}$.
@@ 4. Truncation and reduction
$J=J^{T},\;\;J^{2}=I,$
where $I$ the identity matrix of the same dimension. Now, we invoke the following theorem, which reduces the centrosymmetric matrix to a block diagonal matrix.
@@ 5. Circular Law for Random Centrosymmetric Matrix
Consider a random centrosymmetric matrix $M=\frac{1}{\sqrt{n}}[x_{ij}]_{i,j=1}^{n},$ where $x_{ij}$s are i.i.d. random variables for $i\leq j$ with mean zero and variance $1$ Then, as $n\to\infty$, the ESD of $M$ converges almost surely to the uniform distribution on the unit disc of $\mathbb{C}$ with density $\frac{1}{\pi}1_{\{z\in\mathbb{C}\;|\;|z|\leq 1\}}$.
@@
From Theorem 4.2, it is evident that the eigenvalues of $M$ are the eigenvalues of $T_{1}$ and $T_{2}.$ The entries of $T_{1}$ and $T_{2}$ are i.i.d. centered random variables having variance $2$. However, sizes of $T_{1}$ and $T_{2}$ is $\lceil n/2\rceil\times\lceil n/2\rceil$, and they are scaled by $1/\sqrt{n}.$ Hence, by Theorem 5.1 the ESD of $T_{1}$ converges to the circular law. The same applies to $T_{2}$ as well. Hence, ESD of the matrix $M$ converges to the circular law.
@@ 6. Proof of the central limit theorem
[…]
@@
[…]
where $\hat{\xi}=\xi 1_{\Omega_{n}},$ and $\xi^{\circ}=\xi-\mathbb{E}[\xi].$ The $o(1)$ at the end of the above expression can further be explained as follows.
@@
Next, we prove that $\mathbb{E}[\operatorname{Tr}\mathcal{R}_{z}(M)1_{\Omega_{n}^{c}}]\rightarrow 0$ asymptotically. We note that the assumptions listed in Condition 3.3 are consistent with the assumptions of the Lemma A.3. Hence, using the Lemma A.3, and the bound (6.2) we may conclude that
@@
[…]
@@
where $\sigma$ is a covariance kernel. We defer the calculation of the covariance kernel to Section 7.4. Hence, establishing Proposition 6.1 is equivalent to demonstrating the tightness of the sequence $\{\operatorname{Tr}\hat{\mathcal{R}}_{z}^{\circ}(T_{1})+\operatorname{Tr}\hat{\mathcal{R}}_{z}^{\circ}(T_{2})\}_{n}$ and verifying the conditions outlined in (6.5), (6.6), and (6.7). Before proceeding any further, let us do some reductions.
For $i=1,2$, let $T_{i}^{(k)}$ be the matrix obtained by zeroing out the $k$th column of $T_{i}$, and let ^$\hat{\mathcal{R}}_{z}(T_{i}^{(k)}):=\mathcal{R}_{z}(T_{i}^{(k)})\mathbf{1}_{\Omega_{n,k}}$, where $\Omega_{n,k}=\left\{\left\lVert T_{1}^{(k)}\right\rVert\leq\rho_{1}\right\}\cap\left\{\left\lVert T_{2}^{(k)}\right\rVert\leq\rho_{2}\right\}.$ The space $\Omega_{n,k}$ is different from $\Omega_{n},$ but asymptotically both the spaces are the same. In fact, we can check that
@@
$\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{1})$ $=\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{1}^{(k)})+e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{1}^{(k)})\mathcal{R}_{z}(T_{1})t_{k}^{1}$
$=\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{1}^{(k)})+\frac{e_{k}^{t}(\hat{\mathcal{R}}_{z}(T_{1}^{(k)}))^{2}t_{k}^{1}}{1+e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{1}^{(k)})t_{k}^{1}}.$
@@
$\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{2})=\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{2}^{(k)})+\frac{e_{k}^{t}(\hat{\mathcal{R}}_{z}(T_{2}^{(k)}))^{2}t_{k}^{2}}{1+e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{2}^{(k)})t_{k}^{2}},$
@@
$=\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{1}^{(k)})+\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{2}^{(k)})+\frac{e_{k}^{t}(\hat{\mathcal{R}}_{z}(T_{1}^{(k)}))^{2}t_{k}^{1}}{1+e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{1}^{(k)})t_{k}^{1}}+\frac{e_{k}^{t}(\hat{\mathcal{R}}_{z}(T_{2}^{(k)}))^{2}t_{k}^{2}}{1+e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{2}^{(k)})t_{k}^{2}}$
$=\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{1}^{(k)})+\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{2}^{(k)})-\frac{\partial}{\partial z}\log(1+\gamma_{1k}(z))-\frac{\partial}{\partial z}\log(1+\gamma_{2k}(z)),$
where $\gamma_{1k}(z):=e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{1}^{(k)})t_{k}^{1}$ and $\gamma_{2k}(z):=e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{2}^{(k)})t_{k}^{2}$. Now, let us find the tail bounds of $\gamma_{1k}(z)$ and $\gamma_{2k}(z).$ Note that $\gamma_{1k}(z)$ is product of two independent random variables. Intuitively, $\gamma_{1k}(z)\stackrel{{\scriptstyle a.s.}}{{\to}}0$ by conditioning on $\{t^{1}_{j},t^{2}_{j}:j\in\{1,2,...,\lceil n/2\rceil\}\backslash\{k\}\}$.
Recall that $\sum_{j=1}^{\lceil n/2\rceil}|\hat{\mathcal{R}_{z}}(T_{1}^{(k)})_{ij}|\leq\|\hat{\mathcal{R}_{z}}(T_{1}^{(k)})\|^{2}\leq\tau^{-2}$ for $z\in\partial{\mathbb{D}}_{\rho+\tau}.$ Now, treating $\gamma_{1k}(z)$ as the function $g$ in the property 3 of the Definition 3.2, we obtain
(6.8) $\mathbb{P}\left(|\gamma_{1k}(z)|>t\;\;|\;\;\hat{\mathcal{R}_{z}}(T_{1}^{(k)})\right)\leq K\operatorname{exp}\left(-\tau\sqrt{\frac{\alpha n}{2}}t\right).$
@@
Since $T_{1}^{(k)},T_{2}^{(k)}$ are independent of the $k$th columns of $T_{1}$ and $T_{2}$ respectively, we can easily see that
$\left[\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{1}^{(k)})+\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{2}^{(k)})\right]_{k}^{\circ}$
(6.11) $=\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{1}^{(k)})+\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{2}^{(k)})-\Xi_{k}[\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{1}^{(k)})+\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{2}^{(k)})]=0.$
@@
[…]
@@ 6.1. Proof of tightness
However by Arzelà-Ascoli theorem, the compact sets in space of continuous functions are space of equicontinuous functions. Let us define,
@@
(6.17) $\mathbb{E}\left\{\left|\frac{\operatorname{Tr}\hat{\mathcal{R}}_{z}^{\circ}(T_{1})-\operatorname{Tr}\hat{\mathcal{R}}_{\eta}^{\circ}(T_{1})}{z-\eta}\right|^{2}1_{\Phi_{n}}\right\}\leq C,$
@@
(6.18) $(z-\eta)^{-1}\left[\operatorname{Tr}\hat{\mathcal{R}}_{z}^{\circ}(T_{1})-\operatorname{Tr}\hat{\mathcal{R}}_{\eta}^{\circ}(T_{1})\right]=\sum_{k=1}^{\lceil n/2\rceil}\mathbb{E}_{k}\left\{\operatorname{Tr}\left[\hat{\mathcal{R}}_{k}(T_{1})\hat{\mathcal{R}}_{\eta}(T_{1})\right]_{k}^{\circ}\right\}.$
@@
$\operatorname{Tr}$ [^$\left[\hat{\mathcal{R}}_{z}(T_{1})\hat{\mathcal{R}}_{\eta}(T_{1})-\hat{\mathcal{R}}_{z}\left(T_{1}^{(k)}\right)\hat{\mathcal{R}}_{\eta}\left(T_{1}^{(k)}\right)\right]$
@@
[…]
@@
Thus, using (6.13),(6.14) and the above we have
@@
Expanding $\operatorname{\log}(1+\gamma_{k1}(z))$ and $\operatorname{\log}(1+\gamma_{k2}(z))$ up to two terms and using (6.9),(6.10), and (6.13) we have
@@
$=\frac{1}{n}\left(e_{k}^{t}\mathbb{E}_{k}[\hat{\mathcal{R}}_{z}(T_{1}^{(k)})]\operatorname{Var}((T_{1}^{(k)})_{ki})I\mathbb{E}_{k}[\hat{\mathcal{R}}_{\bar{\eta}}(T_{1}^{(k)}{\operatorname{{}^{*}}})e_{k}]\right)$
$+\frac{1}{n}\left(e_{k}^{t}\mathbb{E}_{k}[\hat{\mathcal{R}}_{z}(T_{2}^{(k)})]\operatorname{Var}((T_{2}^{(k)})_{ki})I\mathbb{E}_{k}[\hat{\mathcal{R}}_{\bar{\eta}}(T_{2}^{(k)}{\operatorname{{}^{*}}})e_{k}]\right)+O(n^{-2})$
@@
Since ^$\|\hat{\mathcal{R}}_{z}(T_{1}^{(k)})+\hat{\mathcal{R}}_{z}(T_{2}^{(k)})\|\leq 4\tau^{-1}$ on $\mathbb{D}_{\rho+\tau/2}^{c}$, the sequence $\chi_{k}(z,\bar{\eta})$ consists of uniformly bounded analytic functions on $\mathbb{D}_{\rho+\tau/2}^{c}$. Thus, by Vitali’s theorem (see [14, Section 2.9]), proving (6.7) is equivalent to show that $\chi_{k}(z,\bar{\eta})$ converges in probability.
@@
$=4e_{k}^{t}\mathbb{E}_{k}[\hat{\mathcal{R}}_{z}(T_{1}^{(k)})]\mathbb{E}_{k}[\hat{\mathcal{R}}_{\bar{\eta}}(T_{1}^{(k)}{\operatorname{{}^{*}}})]e_{k}+4e_{k}^{t}\mathbb{E}_{k}[\hat{\mathcal{R}}_{z}(T_{2}^{(k)})]\mathbb{E}_{k}[\hat{\mathcal{R}}_{\bar{\eta}}(T_{2}^{(k)}{\operatorname{{}^{*}}})]e_{k}$
@@
$\omega_{\epsilon_{n},k}|_{\{\|T_{1}^{(k)}\|\leq\rho_{1}\}}\equiv 1,$
$\omega_{\epsilon_{n},k}|_{\{\|T_{1}^{(k)}\|\geq\rho_{1}+\epsilon_{n}\}}\equiv 0,$
[…]
@@
~$\tilde{\mathcal{R}}_{z}(T_{1}^{(k)})=\mathcal{R}_{z}(T_{1}^{(k)})\omega_{\epsilon_{n},k},$
@@
$\tilde{\Gamma}_{k}^{1}(z,\bar{\eta})=4e_{k}^{t}\mathbb{E}_{k}[\tilde{\mathcal{R}}_{z}(T_{1}^{(k)})]\mathbb{E}_{k}[\tilde{\mathcal{R}}_{\bar{\eta}}(T_{1}^{(k)}{\operatorname{{}^{*}}})]e_{k}.$
Since the smoothed version and the original one match everywhere except the narrow annulus $\{\rho_{1}<\|T_{1}^{(k)}\|<\rho_{1}+\epsilon_{n}\}$, we can easily obtain
$|\Gamma_{k}^{1}(z,\bar{\eta})-\tilde{\Gamma}_{k}^{1}(z,\bar{\eta})|\leq\frac{K}{(\tau-\epsilon_{n})^{2}}1_{\{\rho_{1}<\|T_{1}^{(k)}\|<\rho_{1}+\epsilon_{n}\}},$
@@
Thus, we see that the smoothing function $\omega$ did not alter the original quantities by significant amount. For the notational convenience, let us denote
$\Lambda$ $:=\mathbb{E}_{k}[\mathcal{R}_{z}(T_{1}^{(k)})],$ $\Upsilon$ $:=\mathbb{E}_{k}[\mathcal{R}_{\eta}(T_{2}^{(k)})^{\operatorname{{}^{*}}}],$
$\tilde{\Lambda}$ $:=\mathbb{E}_{k}[\tilde{\mathcal{R}}_{z}(T_{1}^{(k)})],$ $\tilde{\Upsilon}$ $:=\mathbb{E}_{k}[\tilde{\mathcal{R}}_{\eta}(T_{2}^{(k)})^{\operatorname{{}^{*}}}].$

### v3

@@ Abstract.
Indrajit Jana - Email: ijana@iitbbs.ac.in;
Sunita Rani - Email: s21ma09007@iitbbs.ac.in; (Corresponding author).
@@ 2. Notations and layout
This article is organized as follows. We state the main theorem in Section 3. The proof of the CLT is shown in Section 6, and the variance calculation of the limiting normal distribution is delegated to Section 7. In addition, it was also argued that the limiting spectral distribution follows the circular law, which was demonstrated in Sectoin 5.
Throughout this article, the identity matrix is denoted by $I$, and the counter-identity matrix is denoted by $J$, as defined in Definition 4.1. The dimensions of such matrices should be understood from the context. The eigenvalues of an $n\times n$ matrix $M$ are denoted by $\lambda_{1}(M),$ $\lambda_{2}(M),$ $\ldots,$ $\lambda_{n}(M).$ The complex conjugate transpose of a matrix $Y$ is denoted by $Y{\operatorname{{}^{*}}}$, whereas $\bar{Y}$ stands for only the complex conjugate of $Y$ without taking transpose. The notation $\mathcal{R}_{z}(M)=(zI-M)^{-1}$ represents the resolvent of the square matrix $M$. The notations $\zeta_{n}\stackrel{{\scriptstyle d}}{{\to}}\zeta$, $\zeta_{n}\stackrel{{\scriptstyle p}}{{\to}}\zeta$, and $\zeta_{n}\stackrel{{\scriptstyle a.s.}}{{\to}}\zeta$ are used to indicate that the sequence of random variables $\{\zeta_{n}\}$ converges to $\zeta$ in distribution, in probability, and almost surely, respectively.
@@ 3. Main result
Figure 1. Symmetry pattern of a centrosymmetric $\times5$ matrix.
@@
We are now ready to state our main theorem.
@@
$\sigma_{f}^{2}=\frac{2}{\pi}\int_{\mathbb{D}_{1}}f^{\prime}(w)\overline{f^{\prime}(w)}\;d\mathfrak{R}(w)d\mathfrak{I}(w).$
In particular, if $f$ is a polynomial given by $P_{d}(z)=\sum_{k=0}^{d}a_{k}z^{k}$, then $\sigma_{p_{d}}^{2}=\sum_{k=1}^{d}2ka_{k}^{2}$.
@@ 4. Truncation and reduction
$J=J^{T}$
$J^{2}=I,$
where $I$ the identity matrix of the same dimension. Now, we invoke the following theorem which reduces the centrosymmetric matrix to block diagonal matrix.
@@ 5. Circular Law for Random Centrosymmetric Matrix
Consider a random centrosymmetric matrix $M=\frac{1}{\sqrt{n}}[x_{ij}]_{i,j=1}^{n},$ where $x_{ij}$s are i.i.d. random variables for $i\leq j$ with mean zero and variance $1.$ Then, as $n\to\infty$, the ESD of $M$ converges almost surely to the uniform distribution on the unit disc of $\mathbb{C}$ with density $\frac{1}{\pi}1_{\{z\in\mathbb{C}\;|\;|z|\leq 1\}}$.
@@
From Theorem 4.2, it is evident that the eigenvalues of $M$ are the eigenvalues of $T_{1}$ and $T_{2}.$ The entries of $T_{1}$ and $T_{2}$ are i.i.d. centered random variables having variance $2$. However, sizes of $T_{1}$ and $T_{2}$ is $\lceil n/2\rceil\times\lceil n/2\rceil$, and they are scaled by $1/\sqrt{n}.$ Therefore, by Theorem 5.1 the ESD of $T_{1}$ converges to the circular law. The same applies to $T_{2}$ as well. Hence, ESD of the matrix $M$ converges to the circular law.
@@ 6. Proof of the central limit theorem
$=\frac{1}{2\pi i}\oint_{\partial\mathbb{D}_{\rho+\tau}}f(z)\operatorname{Tr}\mathcal{R}_{z}(M)\,dz.$
@@
$\operatorname{L}_{f}^{\circ}(M)=\frac{1}{2\pi i}\oint_{\partial\mathbb{D}_{\rho+\tau}}f(z)\operatorname{Tr}\hat{\mathcal{R}}_{z}^{\circ}(M)\,dz+o(1).$
where $\hat{\xi}=\xi 1_{\Omega_{n}},$ and $\xi^{\circ}=\xi-\mathbb{E}[\xi].$ The $o(1)$ at then end of the above expression can further be explained as follows.
@@
Next, we prove that $\mathbb{E}[\operatorname{Tr}\mathcal{R}_{z}(M)1_{\Omega_{n}^{c}}]\rightarrow 0$ asymptotically.
We note that the assumptions listed in Condition 3.3 are consistent with the assumptions of the Lemma A.3. Therefore, using the Lemma A.3, and the bound (6.2) we may conclude that
@@
$\frac{1}{\pi}\int_{\mathbb{D}_{1}}f^{\prime}(w)\overline{f^{\prime}(w)}\;d\mathfrak{R}(w)d\mathfrak{I}(w)$
$=\frac{1}{\pi}\int_{\mathbb{D}_{1}}\left[\frac{1}{2\pi i}\oint_{\partial\mathbb{D}_{1}}\frac{f(z)\;dz}{(w-z)^{2}}\frac{1}{2\pi i}\oint_{\partial\mathbb{D}_{1}}\frac{f(\bar{\eta})\;d\bar{\eta}}{(\bar{w}-\bar{\eta})^{2}}\right]\;d\mathfrak{R}(w)d\mathfrak{I}(w)$
$=-\frac{1}{4\pi^{2}}\oint_{\partial\mathbb{D}_{1}}\oint_{\partial\mathbb{D}_{1}}f(z)\overline{f(\eta)}\left[\frac{1}{\pi}\int_{\mathbb{D}_{1}}\frac{d\mathfrak{R}(w)d\mathfrak{I}(w)}{(w-z)^{2}(\bar{w}-\bar{\eta})^{2}}\right]\;dz\;d\bar{\eta}$
$=-\frac{1}{4\pi^{2}}\oint_{\partial\mathbb{D}_{1}}\oint_{\partial\mathbb{D}_{1}}\frac{f(z)\overline{f(\eta)}}{(1-z\bar{\eta})^{2}}\;dz\;d\bar{\eta}.$
@@
where $\sigma$ is a covariance kernel to be found later. We defer the calculation of the covariance kernel to section 7.4. Hence, establishing Proposition 6.1 is equivalent to demonstrating the tightness of the sequence $\{\operatorname{Tr}\hat{\mathcal{R}}_{z}^{\circ}(T_{1})+\operatorname{Tr}\hat{\mathcal{R}}_{z}^{\circ}(T_{2})\}_{n}$ and verifying the conditions outlined in (6.5), (6.6), and (6.7). Before proceeding any further, let us do some reductions.
For $i=1,2$, let $T_{i}^{(k)}$ be the matrix obtained by zeroing out the $k$th column of $T_{i}$, and let $\hat{\mathcal{R}}_{z}(T_{i}^{k}):=\mathcal{R}_{z}(T_{i}^{k})\mathbf{1}_{\Omega_{n,k}}$, where $\Omega_{n,k}=\left\{\left\lVert T_{1}^{(k)}\right\rVert\leq\rho_{1}\right\}\cap\left\{\left\lVert T_{2}^{(k)}\right\rVert\leq\rho_{2}\right\}.$ The space $\Omega_{n,k}$ is different from $\Omega_{n},$ but asymptotically both the spaces are the same. In fact, we can check that
@@
$\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{1})$ $=\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{1}^{k})+e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{1}^{k})\mathcal{R}_{z}(T_{1})t_{k}^{1}$
$=\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{1}^{k})+\frac{e_{k}^{t}(\hat{\mathcal{R}}_{z}(T_{1}^{k}))^{2}t_{k}^{1}}{1+e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{1}^{k})t_{k}^{1}}.$
@@
$\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{2})=\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{2}^{k})+\frac{e_{k}^{t}(\hat{\mathcal{R}}_{z}(T_{2}^{k}))^{2}t_{k}^{2}}{1+e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{2}^{k})t_{k}^{2}},$
@@
$=\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{1}^{k})+\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{2}^{k})+\frac{e_{k}^{t}(\hat{\mathcal{R}}_{z}(T_{1}^{k}))^{2}t_{k}^{1}}{1+e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{1}^{k})t_{k}^{1}}+\frac{e_{k}^{t}(\hat{\mathcal{R}}_{z}(T_{2}^{k}))^{2}t_{k}^{2}}{1+e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{2}^{k})t_{k}^{2}}$
$=\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{1}^{k})+\operatorname{Tr}\hat{\mathcal{R}}_{z}(T_{2}^{k})-\frac{\partial}{\partial z}\log(1+\gamma_{1k}(z))-\frac{\partial}{\partial z}\log(1+\gamma_{2k}(z)),$
where $\gamma_{1k}(z):=e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{1}^{k})t_{k}^{1}$ and $\gamma_{2k}(z):=e_{k}^{t}\hat{\mathcal{R}}_{z}(T_{2}^{k})t_{k}^{2}$. Now, let us find the tail bounds of $\gamma_{1k}(z)$ and $\gamma_{2k}(z).$ Note that $\gamma_{1k}(z)$ is product of two independent random variables. Intuitively, $\gamma_{1k}(z)\stackrel{{\scriptstyle a.s.}}{{\to}}0$ by conditioning on $t^{1}_{j},t^{2}_{j}:j\in\{1,2,...,\lceil n/2\rceil\}\backslash\{k\}\}$.
Recall that $\sum_{j=1}^{\lceil n/2\rceil}|\hat{\mathcal{R}_{z}}(T_{1}^{k})_{ij}|\leq\|\hat{\mathcal{R}_{z}}(T_{1}^{k})\|^{2}\leq\tau^{-2}$ for $z\in\partial{\mathbb{D}}_{\rho+\tau}.$ Now, treating $\gamma_{1k}(z)$ as the function $g$ in the property 3 in the Definition 3.2, we obtain
(6.8) $\mathbb{P}\left(|\gamma_{1k}(z)|>t\;\;|\;\;\hat{\mathcal{R}_{z}}(T_{1}^{k})\right)\leq K\operatorname{exp}\left(-\tau\sqrt{\frac{\alpha n}{2}}t\right).$
@@
Since $T_{1}^{(k)},T_{2}^{(k)}$ are independent of the $k$th column of $T_{1},T_{2}$ respectively, we can easily see that
$\left[\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{1}^{k})+\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{2}^{k})\right]_{k}^{\circ}$
(6.11) $=\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{1}^{k})+\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{2}^{k})-\Xi_{k}[\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{1}^{k})+\operatorname{Tr}\hat{\mathcal{R}_{z}}(T_{2}^{k})]=0.$
@@
(6.14) $|v^{\prime}(z)|=\left|\frac{1}{2\pi i}\oint_{z+\partial{\mathbb{D}_{\tau/3}}}\frac{v(t)dt}{(t-z)^{2}}\right|\leq\frac{5}{\pi\tau^{2}}\oint_{z+\partial{\mathbb{D}_{\tau/3}}}|v(t)|dt.$
@@ 6.1. Proof of tightness
However by Arzela-Ascoli theorem, the compact sets in space of continuous functions are space of equicontinuous functions. Let us define,
@@
(6.17) $\mathbb{E}\left\{\left|\frac{|\operatorname{Tr}\hat{\mathcal{R}}_{z}^{\circ}(T_{1})-\operatorname{Tr}\hat{\mathcal{R}}_{\eta}^{\circ}(T_{1})}{z-\eta}\right|^{2}1_{\Phi_{n}}\right\}\leq C,$
@@
(6.18) $(z-\eta)^{-1}\left[\operatorname{Tr}\hat{\mathcal{R}}_{z}^{\circ}(T_{1})-\operatorname{Tr}\hat{\mathcal{R}}_{\eta}^{\circ}(T_{1})\right]=\sum_{k=1}^{\lceil n/2\rceil}\mathbb{E}_{k}\left\{\left(\operatorname{Tr}\left[\hat{\mathcal{R}}_{k}(T_{1})\hat{\mathcal{R}}_{\eta}(T_{1})\right)\right]_{k}^{\circ}\right\}.$
@@
$\operatorname{Tr}$ $\left[\hat{\mathcal{R}}_{z}(T_{1})\hat{\mathcal{R}}_{\eta}(T_{1})-\hat{\mathcal{R}}_{z}\left(T_{1}^{(k)}\right)\hat{\mathcal{R}}_{\eta}\left(T_{1}^{(k)}\right)\right]$
@@
By (6.11), we have $\left[\operatorname{Tr}\left(\hat{\mathcal{R}}_{z}\left(T_{1}^{(k)}\right)\hat{\mathcal{R}}_{\eta}\left(T_{1}^{(k)}\right)\right)\right]_{k}^{\circ}=0$. Therefore, we can express (6.18) as
@@
Thus, using (6.13),(6.14) and the above we have,
@@
Expanding $\operatorname{\log}(1+\gamma_{k1}(z))$ and $\operatorname{\log}(1+\gamma_{k2}(z))$ up to two terms and using (6.9),(6.10) and (6.13) we have
@@
$=\frac{1}{n}\left(e_{k}^{t}\mathbb{E}_{k}[\hat{\mathcal{R}}_{z}(T_{1}^{k})]\operatorname{Var}((T_{1}^{(k)})_{ki})I\mathbb{E}_{k}[\hat{\mathcal{R}}_{\bar{\eta}}(T_{1}^{k}{\operatorname{{}^{*}}})e_{k}]\right)$
$+\frac{1}{n}\left(e_{k}^{t}\mathbb{E}_{k}[\hat{\mathcal{R}}_{z}(T_{2}^{k})]\operatorname{Var}((T_{2}^{(k)})_{ki})I\mathbb{E}_{k}[\hat{\mathcal{R}}_{\bar{\eta}}(T_{2}^{k}{\operatorname{{}^{*}}})e_{k}]\right)+O(n^{-2})$
@@
Since $\|\hat{\mathcal{R}}_{z}(T_{1}^{k})+\hat{\mathcal{R}}_{z}(T_{2}^{k})\|\leq 4\tau^{-1}$ on $\mathbb{D}_{\rho+\tau/2}^{c}$, the sequence $\chi_{k}(z,\bar{\eta})$ consists of uniformly bounded analytic functions on $\mathbb{D}_{\rho+\tau/2}^{c}$. Therefore, by Vitali’s theorem (see [14, Section 2.9]), proving (6.7) is equivalent to show that $\chi_{k}(z,\bar{\eta})$ converges in probability.
@@
$=4e_{k}^{t}\mathbb{E}_{k}[\hat{\mathcal{R}}_{z}(T_{1}^{k})]\mathbb{E}_{k}[\hat{\mathcal{R}}_{\bar{\eta}}(T_{1}^{k}{\operatorname{{}^{*}}})]e_{k}+4e_{k}^{t}\mathbb{E}_{k}[\hat{\mathcal{R}}_{z}(T_{2}^{k})]\mathbb{E}_{k}[\hat{\mathcal{R}}_{\bar{\eta}}(T_{2}^{k}{\operatorname{{}^{*}}})]e_{k}$
@@
$\omega_{\epsilon_{n},k}|_{\{\|T_{1}^{k}\|\leq\rho_{1}\}}\equiv 1,$
$\omega_{\epsilon_{n},k}|_{\{\|T_{1}^{k}\|\geq\rho_{1}+\epsilon_{n}\}}\equiv 0,$
$\left|\frac{\partial}{\partial x_{i}}\omega_{\epsilon_{n},k}\right|\leq\frac{2}{\epsilon_{n}},\;\;\;\;\forall\;\;1\leq i\leq\lceil n^{2}/4\rceil.$
@@
$\tilde{\mathcal{R}}_{z}(T_{1}^{k})=\mathcal{R}_{z}(T_{1}^{k})\omega_{\epsilon_{n},k},$
@@
$\tilde{\Gamma}_{k}^{1}(z,\bar{\eta})=4e_{k}^{t}\mathbb{E}_{k}[\tilde{\mathcal{R}}_{z}(T_{1}^{k})]\mathbb{E}_{k}[\tilde{\mathcal{R}}_{\bar{\eta}}(T_{1}^{k}{\operatorname{{}^{*}}})]e_{k}.$
Since the smoothed version and the original one match everywhere except the narrow annulus $\{\rho_{1}<\|T_{1}^{k}\|<\rho_{1}+\epsilon_{n}\}$, we can easily obtain
$|\Gamma_{k}^{1}(z,\bar{\eta})-\tilde{\Gamma}_{k}^{1}(z,\bar{\eta})|\leq\frac{K}{(\tau-\epsilon_{n})^{2}}1_{\{\rho_{1}<\|T_{1}^{k}\|<\rho_{1}+\epsilon_{n}\}},$
@@
Therefore the smoothing function $\omega$ did not alter the original quantities by significant amount. For the notational convenience, let us denote
$\Lambda$ $:=\mathbb{E}_{k}[\mathcal{R}_{z}(T_{1}^{k})],$ $\Upsilon$ $:=\mathbb{E}_{k}[\mathcal{R}_{\eta}(T_{2}^{k})^{\operatorname{{}^{*}}}],$
$\tilde{\Lambda}$ $:=\mathbb{E}_{k}[\tilde{\mathcal{R}}_{z}(T_{1}^{k})],$ $\tilde{\Upsilon}$ $:=\mathbb{E}_{k}[\tilde{\mathcal{R}}_{\eta}(T_{2}^{k})^{\operatorname{{}^{*}}}].$